# The Negation Of Singer’s Conjecture For The Sixth Algebraic Transfer

### v3

@@ 2 A few preliminaries
It is worth noting that the proof of Theorem 1.4 is presented via our algorithms in SageMath and OSCAR to obviate the need for transcribing unnecessary manual computations. These computations are already detailed in the algorithm’s output (see Note 3.4(B)). Consequently, rather than detailing such lengthy calculations, we sketch the proof based on the construction of our new algorithm.
@@
Detailed computations for this result are shown in the output of our algorithm in Note 3.4(A) below.
@@
$\displaystyle\triangleright$   for all $\mathbf{b}\in$ ExpsEnum($q,n_{g}$do $\triangleright$ stream
@@
    Precomputation of $(\rho_{j}-\mathrm{Id})$ rows on the admissible basis. For each $j=1,\dots,q$ and each admissible monomial $u_{i}$, the row support of $(\rho_{j}-\mathrm{Id})u_{i}$ is computed once and reused across the $\Sigma/GL$ stages and the Stage 4–5 corrections, eliminating repeated decompositions.
@@
where $|\mathbf{Ad}_{6}(\omega_{(i)}|=\dim QP_{6}(\omega^{*}_{(i)})$ for all $i,\,1\leq i\leq 4,$ and the set of all admissible monomials in $\mathbf{Ad}_{6}(\omega_{(i)}$ has also been listed in detail in the output of the algorithm as in Note 3.4(B). Using the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 5,$ and the relation $\rho_{j}(g)\equiv g,$ we see that
@@
where the polynomials $g^{\prime\prime}_{i},\,1\leq i\leq 517,$ are determined from the algorithm output in Note 3.4(B), and
@@
Using this result and Proposition 3.2, we see that if $h\in P_{6}$ such that $[h]\in[(QP_{6})_{36}]^{GL(6)},$ then
@@
Acknowledgment.
We sincerely thank Professor Ulrich Thiel (University of Kaiserslautern-Landau, Germany), an expert in the OSCAR computer algebra system, for the very interesting and useful professional discussions and exchanges. Through these exchanges, we had the opportunity to approach the OSCAR computer algebra system and develop more efficient computational programs.

### v6

@@ 2 A few preliminaries
It is worth noting that the proof of Theorem 1.4 is presented via our algorithms in SageMath and OSCAR to obviate the need for transcribing unnecessary manual computations. These computations are already detailed in the algorithm’s output (see Note 3.5(B)). Consequently, rather than detailing such lengthy calculations, we sketch the proof based on the construction of our new algorithm.
@@
Detailed computations for this result are shown in the output of our algorithm in Note 3.5(A) below.
@@
$\displaystyle\triangleright$   for all $\mathbf{b}\in$ ExpsEnum($q,n_{g}$) do $\triangleright$ stream
@@
$\displaystyle\triangleright$  Output: union of all accepted $GL(q)$-invariants from Step 4 and Step 5 (with logs of weights and dimensions)
@@
    Precomputation of $(\rho_{j}-\mathrm{Id})$ rows on the admissible basis. For each $j=1,\dots,q$ and each admissible monomial $u_{i}$, the row support of $(\rho_{j}-\mathrm{Id})u_{i}$ is computed once and reused across the $\Sigma_{q}/GL(q)$ stages and the Stage 4–5 corrections, eliminating repeated decompositions.
@@
where $|\mathbf{Ad}_{6}(\omega_{(i)}|=\dim QP_{6}(\omega^{*}_{(i)})$ for all $i,\,1\leq i\leq 4,$ and the set of all admissible monomials in $\mathbf{Ad}_{6}(\omega_{(i)}$ has also been listed in detail in the output of the algorithm as in Note 3.5(B). Using the homomorphisms $\rho_{j}:P_{6}\longrightarrow P_{6},\,1\leq j\leq 5,$ and the relation $\rho_{j}(g)\equiv g,$ we see that
@@
where the polynomials $g^{\prime\prime}_{i},\,1\leq i\leq 517,$ are determined from the algorithm output in Note 3.5(B), and
@@
Remark 3.4.
Since ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)})$ is a subspace of $(QP_{6})_{36}$ and the non-zero elements $[\zeta_{1}]$ and $[\zeta_{2}]$ are $GL(6)$-invariants in ${\rm Ker}((\widetilde{Sq}^{0}_{*})_{(6,36)}),$ they are also $GL(6)$-invariants in $(QP_{6})_{36}.$ This implies that
$\dim(\mathbb{F}_{2}\otimes_{GL(6)}\mathcal{P}_{\mathscr{A}}(H_{*}(\mathcal{V}^{6})))_{36}=\dim[(QP_{6})_{36}]^{GL(6)}>\dim{\rm Ext}^{6,6+36}_{\mathscr{A}}(\mathbb{F}_{2},\mathbb{F}_{2})=1.$
Furthermore, our algorithm also finds that $[(QP_{6})_{36}]^{GL(6)}=\mathbb{F}_{2}\cdot([\zeta_{1}],[\zeta_{2}]).$ Indeed, using this result and Proposition 3.2, we see that if $h\in P_{6}$ such that $[h]\in[(QP_{6})_{36}]^{GL(6)},$ then
@@
Funding
D- ặng Võ Phúc was funded by the Post-Doctoral Scholarship Programme of Vingroup Innovation Foundation (VINIF), Institute of Big Data, code: VINIF.2024.STS.38.
Data Availability
The data supporting the findings of this study (specifically, the detailed computational outputs for the cases $(q=5,n=35),$ $(q=6,n=15),$ and $(q=6,n=36)$) are openly available at the URLs provided in the manuscript. The OSCAR source code developed for the computations is available from the corresponding author upon reasonable request.